# Mitigating gain calibration errors from EoR observations with SKA1-Low AA*

### v1

@@ Abstract
The observation of the redshifted 21-cm signal from neutral hydrogen is a promising probe for understanding the phase transition of the early universe and its evolution during the Cosmic Dawn and the Epoch of Reionisation (EoR). One of the primary obstacles to the statistical detection of the target signal is the presence of residual foreground, which arises from gain calibration errors. Previous studies have shown that gain calibration errors as small as 0.01% can lead to a biased interpretation of the observed signal power spectrum estimation, by nearly an order of magnitude. A recent study further highlights that to retrieve astrophysical parameters accurately, the threshold gain calibration error should be lower than 0.01%. In this work, we develop a post-calibration mitigation strategy that combines foreground subtraction techniques with foreground avoidance to address residual contamination arising from gain calibration errors. We use an end-to-end pipeline 21cmE2E to simulate a realistic sky model and telescope configuration within the 138-146 MHz frequency range. To assess the impact of residual antenna-based gain calibration errors, we perform a detailed power spectrum analysis across several threshold levels. Our analysis shows that the extraction of the target signal over the wavenumber range $0.05\leq k\leq 0.5$ Mpc-1 is possible with a threshold gain calibration error of 1%, although with a significant SNR tradeoff on larger scales. Moreover, this work also offers a comparative assessment of foreground removal and avoidance techniques in the context of future SKA1-Low AA* observations.
@@ 1 Introduction
The evolution of the Universe from the post-recombination era to the period following the reionisation of the neutral intergalactic medium (IGM) remains largely unconstrained by direct observations. Theoretical models suggest that the first cosmic structures, comprising the earliest stars and galaxies, formed during this interval, corresponding to redshifts between 30 and 12. The emergence of the first stars, a period referred to as the Cosmic Dawn (CD), was driven by small-scale matter density fluctuations induced by gravitational instabilities. Ionising radiation from these primordial objects gradually ionised the surrounding IGM, initiating the Universe’s final major phase transition. This transformative epoch, known as the Epoch of Reionization (EoR), is believed to span redshifts from approximately 15 to 5 [1, 2].
@@
The neutral hydrogen (Hi) signal from the EoR is expected to be observed at redshifted frequencies in the 50–200 MHz range. Currently, two distinct types of radio telescopes are employed to detect this signal: global signal experiments and interferometers. Global signal experiments use single-radiometer antennas designed to measure the sky-averaged brightness temperature across a broad frequency range. Although EDGES reported an absorption feature centred at 78 MHz [3], subsequent analyses, such as those by SARAS [4], suggest that the best-fit signal may be of non-astrophysical origin. Other single-antenna experiments, including REACH [5] and RHINO [6], are actively pursuing detections of the global signal but are yet to report confirmed results. In parallel, interferometric arrays such as LOFAR [^2] Low Frequency Array, http://www.lofar.org [7], MWA [^3] Murchison Widefield Array, http://www.mwatelescope.org [8], HERA [^4] Hydrogen Epoch of Reionization Array, http://reionization.org [9], uGMRT [^5] Giant Metrewave Radio Telescope , https://www.gmrt.ncra.tifr.res.in/ [10], and the forthcoming SKA [^6] Square Kilometre Array, https://www.skao.int/en [11] aim to detect spatial fluctuations in differential brightness temperature statistically. Although these experiments have placed upper limits on the EoR power spectrum [12, 13, 14, 15], a definitive detection of the Hi signal remains elusive.
@@
The 21-cm signal is susceptible to contamination by bright astrophysical foregrounds, primarily galactic diffuse synchrotron and free-free emissions, along with extragalactic radio sources such as star-forming galaxies and Active Galactic Nuclei (AGN) [16, 17, 18, 19, 20]. These foregrounds are expected to be 4–5 orders of magnitude brighter than the Hi signal. Mitigation of these foregrounds is particularly challenging for interferometric experiments, as global signal experiments benefit from larger beam areas that suppress contributions from bright extragalactic point sources [21]. However, foregrounds in both approaches typically follow a power-law spectrum, leading to inherent spectral smoothness. Since the Hi signal arises against the cosmic microwave background and propagates through an isotropically expanding Universe, its spherical symmetry is preserved in Fourier space, allowing the construction of isotropic image cubes [22]. Unlike foregrounds, the Hi signal exhibits discrete frequency structures due to redshift–frequency mapping, corresponding to line-of-sight (LoS) distances. These contrasting spectral properties form the basis for foreground mitigation strategies.
@@
The effectiveness of foreground mitigation is typically evaluated using cylindrically averaged power spectra (2D PS), which distinctly separate the EoR signal, astrophysical foregrounds, and instrumental systematics in two-dimensional ($k_{\perp},k_{\parallel}$) $k$-space. Smooth foregrounds are expected to dominate low $k_{\parallel}$ modes, contributing mostly around zero delay, while the Hi signal is confined to higher $k_{\parallel}$ values, forming the so-called “EoR window” [23]. The limits in $k_{\perp}$ and $k_{\parallel}$ are governed by the baseline distribution of the array and spectral bandwidth/resolution, respectively. However, instrument chromaticity leads to mode mixing, where foreground power leaks into the EoR window, forming the “foreground wedge” [24, 25, 26]. The statistical detection of the Hi signal is performed using the spherically averaged power spectrum (1D PS).
@@
Foreground mitigation approaches are broadly classified into foreground removal and foreground avoidance [24, 27]. Although both of these approaches have contributed to current upper limits, each has tradeoffs, and no consensus exists on a preferred approach for SKA-Low. Foreground avoidance is robust to systematics but sacrifices sensitivity, especially on large scales, as it discards low $k_{\parallel}$ modes [28]. Conversely, non-parametric subtraction techniques [29, 30, 31, 32, 33, 34], which rely on the frequency-frequency covariance of statistically independent components, are susceptible to signal loss due to leakage of the cosmological signal into the foreground model. The situation is worsened by systematics such as calibration errors, incomplete sky models [35, 36, 37, 38, 39, 40, 41], beam uncertainties [42, 43, 44], mutual coupling [45, 46, 47], and ionospheric distortions [48, 49]. These effects can introduce excess variance or mimic the spectral characteristics of the Hi signal, leading to false detections.
@@
In this work, we focus on mitigating the impact of gain calibration errors using post-calibration residual foreground mitigation techniques. We assess the efficacy of Principal Component Analysis (PCA) and Gaussian Process Regression (GPR) in recovering the 21 cm power spectrum under varying gain error conditions. In addition to purely foreground avoidance or subtraction, inspired by ’hybrid’ foreground mitigation techniques that suggest avoidance along with subtraction [50, 51, 52], we apply PCA and GPR to suppress foreground power at low $k_{\parallel}$ sufficiently to open the EoR window for avoidance-based analysis. We demonstrate that this hybrid approach allows partial recovery of large-scale $k$-modes in the 2D PS, where the Hi signal is strongest, and hence improves the signal-to-noise ratio (SNR). Our primary objective is to mitigate the residual foreground resulting from gain calibration errors as much as possible, while maintaining power spectrum sensitivity.
@@ 2 Simulations
This section discusses the simulation framework for introducing residual gain calibration errors into SKA-Low AA* data through synthetic radio observations using end-to-end simulations.
@@ 2.1 Synthetic observations
We used a 21-cm end-to-end (21cmE2E) pipeline [39, 48, 54] to perform synthetic observations. This pipeline employs the OSKAR [^7] https://github.com/OxfordSKA/OSKAR software package [55] for simulating visibilities for SKA1-Low configurations. In this simulated observation, the sky was observed with a phase-center at $\alpha$ = 15h00m00s and $\delta$ = -30 for a duration of 4 hours ($\pm$ 2 HA). The observing bandwidth of the lightcone spans 8 MHz with a channel separation of 125 kHz. The observation parameters are listed in the table 1. In the subsection, we describe the realistic sky models and telescope configuration to perform the simulation using the 21cmE2E pipeline.
@@
2.1.1 EoR Signal
For Hi simulations, we use the semi-numerical EoR modelling tool, 21cmFAST[56, 57], to generate the brightness temperature map ($\delta T_{b}$). We consider a spatial comoving volume of $500\times 500\,$h-1Mpc2 and a redshift range of 7.4 to 8.2 with pixel size of $232\times 232$ and channel width of 125 kHz. This work used the existing simulated Hi 21-cmmaps from Tripathi et al. [58]. The left plot of Figure 2 shows a 3x3 slice of Hi map at 144.25 MHz. The top-left plot of Figure 3 shows the 2D PS, and the red markers in Figure 4 show the 1D PS of one such realisation of the Hi cube as observed with our telescope model. Since we do not include thermal noise in our data, we consider this to be the ground truth for all power spectrum results.
2.1.2 Point source model
The foreground model used in this study consists exclusively of compact sources extracted from the Tiered Radio Extragalactic Continuum Simulation (T-RECS) catalogue [59]. The sources were redistributed around a declination of -30  and the flux densities were interpolated to 142 MHz by assuming a smooth power law relation with spectral index $\alpha$ equal to -0.8. Our sky model consists of 2522 point sources within our simulated FoV of 4. We refer to Section 2 of [48] for a comprehensive overview of the sky model. It is important to note that the diffuse foreground models are not included in our simulation.
@@ 2.1.3 Telescope model
The SKA1-Low AA* [60] layout is a defined configuration of the SKA1-Low array, comprising 307 stations with a maximum baseline of approximately 73.4 km [61]. It preserves the essential features of the full array design, with a dense central core, where roughly 50% of the stations are packed within a 1 km diameter, and an extended distribution along three logarithmically spaced spiral arms [62]. This layout balances the sensitivity to surface brightness and the angular resolution, crucial for tomographic imaging and statistical studies of EoR [63]. In our work, we restrict the array to stations within a 2 km radius to enhance the sampling of short baselines and thereby improve sensitivity to the cosmologically relevant large-scale 21-cm fluctuations while reducing computational demands. The telescope configuration of SKA1-Low AA* is shown in Figure 1 and the corresponding baseline coverage in the UV plane for an observation time of 4 hours ($\pm 2$ HA).
@@ 2.1.4 Gain Calibration Errors
Gain calibration errors are defined as imperfections in the amplitude and phase of these complex gain terms, which may arise from various instrumental factors, including temperature fluctuations in electronic components, mutual couplings, and time-dependent variations in antenna response. For the scope of this work, we have excluded calibration bandpass errors, assuming that frequency-dependent effects are perfectly corrected and thus do not contribute to the residual errors in our analysis.
The errors due to inaccurate calibration can introduce spurious features that either mimic or entirely obscure the faint 21-cmcosmological signal [40]. Simulations of frequency-independent calibration errors by Morales et al. [64] suggest a more prominent contamination in the reconstructed (image-based) power spectrum than in the delay spectrum estimates. Therefore, we expect that the contamination in our estimation also depends on the quality of the instrumental Point Spread Function (PSF), as pointed out by [35]. Mazumder et al. [39] showed that for SKA1-Low, gain calibration errors as low as 0.1% can significantly contaminate the EoR window, thereby reducing the signal-to-noise ratio. The primary goal of this work is to mitigate the effects of these calibration errors during post-processing, treating them as perturbations in the gain solutions that arise from an unoptimised calibration algorithm.
@@ 2.1.5 Imaging
The synthesised maps were generated from the simulated visibilities from the 21cmE2E pipeline using WSCLEAN [65]. The visibilities were gridded, and each sub-band was imaged independently. We stack each sub-band along the frequency axis, making an image cube ($\theta_{x},\theta_{y},\nu$). We apply Gaussian tapering and W-projection to produce naturally weighted dirty images of 4 FoV with a pixel size of $15\times 15$  arcsec. In figure 2, we show one such slice of the image cube of the observed Hi signal maps via SKA-Low simulations. Although we simulate 4 sky areas around the zenith for all the components of the sky signal, we perform the power spectrum estimation using only 3 from the phase-centre, corresponding to SKA-low’s anticipated primary beam [66]. The maximum and minimum $k_{\perp}$ corresponding to image size and PSF are 0.036 Mpc-1 to 0.63 Mpc-1. We find that the effect of PSF starts dominating around $k_{\perp}$ $\sim$ 0.4 Mpc-1. Out of the spherically averaged $k$-modes probable by our observations 0.05 $\leq$ k $\leq$ 1.57 Mpc-1, we focus on large scales of cosmological interest and perform power spectrum analysis for 0.05 $\leq$ k $\leq$ 0.5 Mpc-1.
@@ 2.2 Power spectrum estimation
where ‘i’ corresponds to gain error, PCA or GPR-filtered power spectrum.
@@
Figure 3: Proceeding clockwise from top-left, we show cylindrically averaged 2-D PS for the true Hi (perfect foreground subtraction), and the reconstructed PS of the residual data with 0.1%, 1%, and 10% gain calibration error. The minimum and maximum baselines were taken to be $\sim 48\lambda$ and $\sim 948\lambda$ and the black and white dotted lines represent the horizon limit and horizon plus 0.1 Mpc-1 buffer, respectively.
@@
$\sigma^{2}_{i}=\frac{\Delta^{2}(k\in k_{i})}{\sqrt{N_{i}}}$ (2.11)
@@
where $\Delta_{\rm HI}^{2}(k)$ is the Hi power spectrum and $\Delta_{N}^{2}(k)$ is the total uncertainty (sample variance + 100 hr of observation of instrumental noise).
As can be seen from the 2-D cylindrical power spectrum in Figure 3, residual foreground contamination in the EoR window increases with increasing residual gain calibration error. Specifically, for the calibration inaccuracy of 0.1%, the excess residual power manifests itself as a mode-mixing feature [34] at high $k_{\perp}$ modes. Thus, foreground avoidance about the horizon line appears viable (see Section 4.1). However, for higher corruption levels, the familiar foreground wedge, which includes both intrinsic smooth partitions and mode-mixing features, becomes prominent.

### v2

@@ Abstract
[…]
@@ 1 Introduction
The evolution of the Universe from the post-recombination era to the period following the reionisation of the neutral intergalactic medium (IGM) remains largely unconstrained by direct observations. Theoretical models suggest that the first cosmic structures, comprising the earliest stars and galaxies, formed during this interval, corresponding to redshifts between 30 and 7 [1]. The emergence of the first stars, a period referred to as the Cosmic Dawn (CD), was driven by small-scale matter density fluctuations induced by gravitational instabilities. Ultraviolet (UV) radiation from these primordial objects gradually ionised the surrounding IGM, initiating the Epoch of Reionization (EoR) [2, 3].
@@
The neutral hydrogen (Hi) signal from the EoR is expected to be observed at redshifted frequencies in the 50–200 MHz range. Currently, two distinct approaches of observations are employed to detect this signal. Global signal experiments consisting of single radiometer antennas aim to detect the fluctuation of sky brightness temperature across a broad frequency range, whereas radio interferometers aim to detect spatial fluctuations in sky brightness temperature across redshifts. Although EDGES reported an absorption feature centred at 78 MHz [4], subsequent analyses, such as those by SARAS [5], suggest that the best-fit signal may be of non-astrophysical origin. Other single-antenna experiments, including REACH [6] and RHINO [7], are actively pursuing detections of the global signal but are yet to report confirmed results. Interferometric arrays such as LOFAR [^2] Low Frequency Array, http://www.lofar.org [8], MWA [^3] Murchison Widefield Array, http://www.mwatelescope.org [9], HERA [^4] Hydrogen Epoch of Reionization Array, http://reionization.org [10], uGMRT [^5] Giant Metrewave Radio Telescope , https://www.gmrt.ncra.tifr.res.in/ [11], and the forthcoming SKA [^6] Square Kilometre Array, https://www.skao.int/en [12] aim to detect spatial fluctuations in differential brightness temperature statistically. Although these experiments have placed upper limits on the EoR power spectrum [13, 14, 15, 16], a definitive detection of the Hi signal remains elusive.
@@
One of the major observational challenges to the detection of the Hi signal is bright astrophysical foregrounds, primarily galactic diffuse synchrotron and free-free emissions, along with extragalactic radio sources such as star-forming galaxies and Active Galactic Nuclei (AGN) [17, 18, 19, 20, 21]. These foregrounds are expected to be 4–5 orders of magnitude brighter than the Hi signal. However, the foregrounds typically follow a power-law spectrum, leading to inherent spectral smoothness. In contrast, the Hi signal exhibits spectral structures since each frequency corresponds to a different redshift along the line-of-sight (LoS) distances. These contrasting spectral properties form the basis for foreground mitigation strategies.
@@
The effectiveness of foreground mitigation is typically evaluated using cylindrically averaged angular two-dimensional power spectra (2D PS) which contain distinct regions where the EoR signal, astrophysical foregrounds, and instrumental systematics dominate. The y-axis of 2D PS represents the Fourier transformation along the frequency direction ($k_{\parallel}$) to the delay time space for individual baselines, a technique called the ‘delay transform’ [22]. The x-axis represents the Fourier transformation of angular scales, perpendicular to the LoS distances ($k_{\perp}$). The limits on $k_{\perp}$ and $k_{\parallel}$ are governed by the baseline distribution of the array and spectral bandwidth/resolution, respectively. Spectrally smooth foregrounds are expected to dominate low $k_{\parallel}$ modes, contributing mostly around zero delay. While the Hi signal is strongest on larger scales (low $k_{\perp}$), it remains uncontaminated towards higher $k_{\parallel}$ values, dominated only by thermal noise, forming the so-called “EoR window” [23]. However, instrument chromaticity leads to mode mixing, where foreground power leaks into the EoR window, forming the “foreground wedge” [24, 25, 26].
@@
Foreground mitigation approaches are broadly classified into foreground removal and foreground avoidance [24, 27]. Foreground avoidance is an approach that involves discarding the contaminated modes inside the foreground wedge, probing the accessible EoR window. On the other hand, foreground removal techniques aim to fit a foreground model and subtract it, preserving power spectrum sensitivity across all scales. Although both of these approaches have contributed to current upper limits, each has tradeoffs. Foreground avoidance is robust to systematics but sacrifices sensitivity, especially at large scales, because it discards low $k_{\parallel}$ modes [28]. Conversely, non-parametric subtraction techniques [29, 30, 31, 32, 33, 34], which rely on the frequency-frequency covariance of statistically independent components, are susceptible to signal loss due to leakage of the cosmological signal into the foreground model [35]. The inaccuracies in the foreground model are enhanced by systematics such as calibration errors, incomplete sky models [36, 37, 38, 39, 40, 41, 42], beam uncertainties [43, 44, 45], mutual coupling [46, 47, 48], and ionospheric distortions [49, 50]. These effects can introduce excess variance or mimic the spectral characteristics of the Hi signal, leading to false detections.
@@
In this work, we focus on mitigating the impact of gain calibration errors using post-calibration residual foreground mitigation techniques. We use an end-to-end simulations pipeline to assess the efficacy of Principal Component Analysis (PCA) [33] and Gaussian Process Regression (GPR) [34] in recovering the 21 cm power spectrum under varying gain error conditions. In addition to purely foreground avoidance or subtraction, inspired by ’hybrid’ foreground mitigation techniques that suggest avoidance along with subtraction [51, 52, 53], we apply PCA and GPR to suppress foreground power at low $k_{\parallel}$ sufficiently to open the EoR window for avoidance-based analysis. We demonstrate that this hybrid approach allows partial recovery of large-scale $k$-modes in the 2D PS, where the Hi signal is strongest, and hence improves the signal-to-noise ratio (SNR). Our primary objective is to mitigate the residual foreground resulting from gain calibration errors as much as possible, while maintaining power spectrum sensitivity.
@@ 2 Simulations
In this section, we present an end-to-end framework that simulates synthetic SKA-Low AA* data, introduces residual gain calibration errors, and produces corrupted visibilities as a result.
@@ 2.1 Synthetic observations
Foregrounds (FG)Point SourcesHi Signal(21cmFAST)Telescope LayoutSKA1-LOW AA*OSKAR SimulationsModel VisibilityCalibration Gain errorResidual Visibility(Model-Corrupted)Foreground MitigationSubtraction(PCA,GPR)Avoidance(horizon or horizon + buffer)Hybrid Methods(subtraction+avoidance)Power Spectrum1D,2D PS
Figure 1: A flowchart depicting the workflow of 21cmE2E, along with the residual foreground mitigation techniques and power spectra analysis done in this work.
We used a 21-cm end-to-end (21cmE2E) pipeline [40, 49, 55] to perform synthetic observations for the SKA-Low telescope layout AA*. This pipeline employs the OSKAR [^7] https://github.com/OxfordSKA/OSKAR software package [56] for simulating visibilities for SKA1-Low configurations. The resulting measurement sets were then processed using the Common Astronomy Software Applications (CASA) [57] package for reading, calibration, and further analysis of the visibility data. In this simulated observation, the sky was observed with a phase-center at $\alpha$ = 15h00m00s and $\delta$ = -30 for a duration of 4 hours ($\pm$ 2 HA). The observing bandwidth of the lightcone spans 8 MHz with a channel separation of 125 kHz. A flow chart showing the workflow of 21cmE2E pipeline integrated by foreground mitigation techniques is shown in the Figure 1. The observation parameters used in the simulations are listed in the Table 1. In the subsections, we describe the realistic sky models and telescope configuration used to perform the simulation with the 21cmE2E pipeline.
@@ 2.1.1 Telescope model
The SKA1-Low AA* [58] layout (Figure 2) is a defined configuration of the SKA1-Low array, comprising 307 stations (at the time of writing) with a maximum baseline of approximately 73.4 km [59]. It preserves the essential features of the full array design, with a dense central core, in which roughly 50% of the stations are packed within a 1 km diameter, and an extended distribution along three logarithmically spaced spiral arms [60]. In our work, we restrict the array to stations within a 2 km radius to mitigate foreground leakage on larger scales [61], also reducing computational demands.
@@
2.1.2 EoR Signal
For Hi simulations, we use the semi-numerical EoR simulation, 21cmFAST[62, 63], to generate the brightness temperature map ($\delta T_{b}$). We consider a spatial co-moving volume of $500\times 500\,$h-1Mpc2, with pixel size of $232\times 232$, a redshift range of 7.4 to 8.2 and channel width of 125 kHz respectively. This work used the existing simulated Hi lightcone from Tripathi et al. [64]. Since we do not include thermal noise in our data, we consider this to be the ground truth for all power spectrum results.
2.1.3 Point source model
The foreground model used in this study consists exclusively of compact sources extracted from the Tiered Radio Extragalactic Continuum Simulation (T-RECS) catalogue [65]. The sources were redistributed around a declination of -30  and the flux densities were interpolated to 142 MHz by assuming a smooth power law relation with spectral index $\alpha$ equal to -0.8. Our sky model consists of 2522 point sources within our simulated FoV of 4. We refer to Section 2 of [49] for a comprehensive overview of the sky model. It is important to note that the galactic diffuse foreground models are not included in our simulation.
@@ 2.1.4 Gain Calibration Errors
Gain calibration errors are defined as imperfections in the amplitude and phase of these complex gain terms, which may arise from various instrumental factors, including temperature fluctuations in electronic components, and time-dependent variations in antenna response [36]. For the scope of this work, we have excluded calibration bandpass errors.
The errors due to imperfect source model calibration can introduce spurious features that either mimic or entirely obscure the faint 21-cm cosmological signal [41]. Simulations of frequency-independent calibration errors by Morales et al. [66] suggest a more prominent contamination in the reconstructed (image-based) power spectrum than in the visibility-based power spectrum, estimates also known as the delay spectrum. Therefore, we expect that the contamination in our estimation also depends on the quality of the instrumental Point Spread Function (PSF) [36]. Mazumder et al. [40] showed that for SKA1-Low, gain calibration errors as low as 0.01% can significantly contaminate the EoR window, thereby reducing the signal-to-noise ratio. The primary goal of this work is to mitigate the effects of these calibration errors during post-processing, treating them as perturbations in the gain solutions that arise from an unoptimised calibration algorithm.
@@ 2.1.5 Imaging
The synthesised maps were generated from the simulated visibilities from the 21cmE2E pipeline using WSCLEAN [67]. The visibilities were gridded, and each sub-band was imaged independently. We stack each sub-band along the frequency axis, making an image cube ($\theta_{x},\theta_{y},\nu$). We apply Gaussian tapering and W-projection to produce naturally weighted dirty images of 4 FoV with a pixel size of $15\times 15$  arcsec. The plots at the bottom of Figure 3 show a 3x3 slice of Hi map and point source subtracted residual map with 1% at 144.25 MHz, respectively. Although we simulate 4 sky areas around the zenith for all the components of the sky signal, we perform the power spectrum estimation using only 3 from the phase-centre, corresponding to SKA-low’s anticipated primary beam [68]. The effect of PSF starts dominating around $k_{\perp}$ $\sim$ 0.38 Mpc-1. Out of the spherically averaged $k$-modes available by our observations 0.05 $\leq$ k $\leq$ 1.57 Mpc-1, we focus on large scales of cosmological interest and perform power spectrum analysis for 0.05 $\leq$ k $\leq$ 0.5 Mpc-1.
@@ 2.2 Power spectrum estimation
where ‘i’ corresponds to gain error, PCA-, or GPR-filtered power spectrum.
@@
Figure 4: From top to bottom, we show cylindrically averaged 2-D PS for three scenarios: no FG avoidance, avoidance about the horizon, and avoidance about horizon plus 0.1 Mpc-1 buffer, respectively. Proceeding right to left, we show 2-D PS for the true Hi (perfect foreground subtraction), and for the point-source subtracted residual data with 0.1%, 1%, and 10% gain calibration error. The minimum and maximum baselines were taken to be $\sim 48\lambda$ and $\sim 948\lambda$. The black and white dotted lines represent the horizon limit and horizon plus 0.1 Mpc-1 buffer, respectively. The vertical red dotted line marks the scale at which PSF starts to dominate.
@@
$\sigma^{2}_{i}=\frac{\Delta^{2}(k\in k_{i})}{N_{i}}$ (2.11)
@@
where $\Delta_{\rm HI}^{2}(k)$ is the Hi power spectrum and $\Delta_{N}^{2}(k)$ is the total uncertainty (sample variance + 100 hrs of observation of instrumental noise).
The plots in the last column of Figure 4, and the gray markers in Figure 5 show the 2D PS and 1D PS of a realisation of the Hi cube as observed with our telescope model. As can be seen from the 2-D cylindrical power spectrum in Figure 4, residual foreground contamination in the EoR window increases with increasing residual gain calibration error. Specifically, for the calibration inaccuracy of 0.1%, the excess residual power manifests itself as a mode-mixing feature [34] at high $k_{\perp}$ modes. Thus, foreground avoidance about the horizon line appears viable (see Section 4.1). However, for higher corruption levels, the familiar foreground wedge, which includes both intrinsic smooth partitions and mode-mixing features, becomes prominent.